OpenStudy (anonymous):

31. Hiram raises earthworms. In a square of compost 4 ft by 4 ft, he can have 1,000 earthworms. How many earthworms can he have if his square of compost has a side length that is 5 times longer? (1 point)

OpenStudy (anonymous):

find the area= 4 * 4 = 16 ft^2 and if the side is 5 times longer then, 4(5) * 4(5) = 400ft^2 Use ratio if 16ft^2 = 1000 worms then 400ft^2 = x worms? 1000*400/16 = x